3. A building with an appraisal value of $135,757 is made available at an offer price of $158,563. The purchaser acquires the property for $34,867 in cash, a 90-day note payable for $23,496, and a mortgage amounting to $58,384. The cost basis recorded in the buyer's accounting records to recognize this purchase is

a.$158,563

b.$135,757

c.$123,696

d.$116,747

4. A machine with a cost of $135,000 has an estimated residual value of $15,000 and an estimated life of 6 years or 15,000 hours. It is to be depreciated by the units-of-activity method. What is the amount of depreciation for the second full year, during which the machine was used 5,100 hours?

a.$51,000

b.$40,800

c.$25,800

d.$5,100

5. On December 31, Strike Company has decided to discard one of its batting cages. The equipment had an initial cost of $205,000 and has accumulated depreciation of $184,500. Depreciation has been recorded up to the end of the year. Which of the following will be included in the entry to record the disposal?

a.Equipment, credit, $205,000

b.Gain on Disposal of Asset, credit, $20,500

c.Loss on Disposal of Asset, debit, $184,500

d.Accumulated Depreciation, debit, $205,000

6. Equipment with a cost of $154,821 has an estimated residual value of $8,049 and an estimated life of 4 years or 15,860 hours. It is to be depreciated by the straight-line method. What is the amount of depreciation for the first full year, during which the equipment was used for 3,136 hours?

a.$36,693.00

b.$29,021.25

c.$38,705.25

d.$29,068.05

7.

Computer equipment was acquired at the beginning of the year at a cost of $58,800. It had an estimated residual value of $4,800 and an estimated useful life of five years. Determine the (a) depreciable cost, (b) straight-line rate, and (c) annual straight-line depreciation.

a. Depreciable cost $
b. Straight-line rate %
c. Annual straight-line depreciation $
Answer #1

3.

Cost basis recorded in buyer's accounting records = $34,867+23,496+58,384 = $116,747

Hence option d is correct.

4.

Depreciation for second year = $135,000 - 15,000 / 15,000 * 5,100 = $40,800

Hence option b is correct.

5.

Journal entry to record disposal

General Journal Debit Cedit
Accumulated depreciation $184,500
Loss on disposal of asset 20,500
Equipment $205,000

Hence option a is correct

6.

Depreciation for full year = $154,821 - 8,049 / 4 = $36,693

Hence option a is correct

7.

a. Depreciable cost = $58,800 - 4,800 = $54,000

b. Straight line rate = 100 / 5 years = 20%

c. Annual straight line depreciation = $10,800